# Analysis of Broken Randomized Experiments by Principal Stratification

### v1

@@ 2.2 Identification
$P(S(0)=1\mid c)$ =$=\frac{\{p_{c}+p_{n}\}E\{S\mid Z=0,D=0\}-p_{n}E\{S\mid Z=1,D=0\}}{p_{c}}$
The third-stage regression links the outomes with principal strata. Note that $YS$ is always well defined, so we use $YS$ as the outcome.
@@
$E\{Y(1)\mid G=al\}$ $=\frac{E\{YS\mid Z=0,D=1\}}{P(S(1)=1\mid a)}$
$=\frac{E\{S\mid(1-Z)D\Delta_{S}=1\}E\{Y\mid(1-Z)DS\Delta_{Y}=1\}}{E\{S\mid Z=0,D=1\}},$
$E\{Y(0)\mid G=nl\}$ $=\frac{E\{YS\mid Z=1,D=0\}}{P(S(0)=1\mid n)}$
$=\frac{E\{S\mid Z(1-D)\Delta_{S}=1\}E\{Y\mid Z(1-D)S\Delta_{Y}=1\}}{E\{S\mid Z=1,D=0\}}.$
Afterwards, the identifiability result follows immediately.
@@
When the way of breaking randomized experiments is simple, we can simplify the identification results.
@@
In the presence of missingness, we can simply use imputation methods to impute the missing values, and use the formula in Remark 3 to estimate the principal average causal effect (Harel and Zhou, , 2007; Kenward and Carpenter, , 2007; Rubin, , 2018). To be noted, the proportions of principal strata $cl$, $cp$, $ch$, $cd$ defined by compliance type and survival type are not identifiable without additional assumptions. A typical assumption is the individual-level monotonicity on $S$, i.e., $S(1)\geq S(0)$ almost surely, under which
@@
Whether this assumption is plausible depends on specific scientific questions. For example, this monotonicity does not hold for JC data (Frumento et al., , 2012) but may be reasonable in some medical applications (Deng et al., , 2021).
The unidentifiability of principal strata proportions brings difficulty in aggregating the conditional causal effect in each covariates level to the average causal effect in the overall population. In the presence of baseline covariates $X$, we can impose the identification assumptions in each level of $X$ and identify the conditional principal average causal effect (CPACE)
$\tau(x)=E\{Y(1)-Y(0)\mid G=cl,X=x\}.$
However, since we do not know $P(G=cl\mid X=x)$, we can not aggregate the condition effects
[…]
Therefore, the principal stratification analysis on broken randomized experiments should be conducted in each level of covariates. Otherwise we need to impose additional assumptions; see Appendix for more details.
@@ 3 Estimation and inference
Consider a completely randomized design with known treatment propensity $P(Z=1)=1-P(Z=0)=r$. Suppose there is no missingness, otherwise one can use multiple imputation or other methods to generate a completed dataset first. The asymptotic variance of the estimator can be calculated by Rubin’s rule from completed datasets (Rubin and Schenker, , 1986; White et al., , 2011; Rubin, , 2018).
@@
$\theta=(\theta_{1},\theta_{0},\theta_{11},\theta_{10},\theta_{01},\theta_{00},\theta_{111},\theta_{101},\theta_{011},\theta_{001})^{{\mathrm{\scriptscriptstyle T}}}.$
The likelihood with respect to $\theta$ is
@@
Solving the maximum likelihood estimation equation $\partial\log L(\theta;\mathcal{O})/\partial\theta=0$, one can obtain
$\widehat{\theta}=\left(\frac{n_{11\cdot}}{n_{1\cdot\cdot}},\frac{n_{01\cdot}}{n_{0\cdot\cdot}},\frac{n_{111}}{n_{11\cdot}},\frac{n_{101}}{n_{10\cdot}},\frac{n_{011}}{n_{01\cdot}},\frac{n_{001}}{n_{00\cdot}},\widehat{\theta}_{111},\widehat{\theta}_{101},\widehat{\theta}_{011},\widehat{\theta}_{001}\right)^{{\mathrm{\scriptscriptstyle T}}},$
@@
$\operatorname{cov}(\widehat{\theta})$ $=\operatorname{diag}\bigg{\{}\frac{r}{n\theta_{1}(1-\theta_{1})},\frac{1-r}{n\theta_{0}(1-\theta_{0})},\frac{r\theta_{1}}{n\theta_{11}(1-\theta_{11})},\frac{r(1-\theta_{1})}{n\theta_{10}(1-\theta_{10})},$
$\qquad\qquad\frac{(1-r)\theta_{0}}{n\theta_{01}(1-\theta_{01})},\frac{(1-r)(1-\theta_{0})}{n\theta_{00}(1-\theta_{00})},\operatorname{var}(\widehat{\theta}_{111}),\operatorname{var}(\widehat{\theta}_{101}),\operatorname{var}(\widehat{\theta}_{011}),\operatorname{var}(\widehat{\theta}_{001})\bigg{\}}.$
If we use the sample average $\widehat{\theta}_{zd1}$ of the outcomes in $\{Z=z,D=d,S=1\}$ to estimate $b(\theta_{zd1})$, then the variance $\operatorname{var}(\widehat{\theta}_{zd1})$ is the standard deviation of the outcomes in $\{Z=z,D=d,S=1\}$ divided by the group size $n_{zd1}$.
Since we can express
@@
the causal paratemers can be estimated by plugging $\widehat{\theta}$ into $\mu_{d}$, denoted by $\widehat{\mu}_{d}$, for $d=0,1$. Similarly, the principal average causal effect is estimated by $\widehat{\tau}=\widehat{\mu}_{1}-\widehat{\mu}_{0}$. The asymptotic properties of the plug-in estimators can be established by sandwich estimator, as shown in the following theorem.
@@
When the outcome $Y$ is binary, the normal approximation may not be good, since the distribution of the estimator $\widehat{\mu}_{d}$ can be skewed. To respect the fact that $\mu_{d}$ must lie in $[0,1]$, we may as well adopt $\operatorname{logit}(\mu_{d})=\log\{\mu_{d}/(1-\mu_{d})\}$ as the target estimand, which can range on the whole real line. The contrast $\tau^{*}=\operatorname{logit}(\mu_{1})-\operatorname{logit}(\mu_{0})$ reflects the log odds ratio between treatment and control for survived compliers. However, the log odds ratio does not enjoy collapsibility. Since $Y$ is binary, we have $\theta_{zd1}=E\{Y\mid Z=z,D=d,S=1\}$. Now
$\operatorname{logit}(\mu_{1})$ $=\log\left\{\frac{\theta_{1}\theta_{11}\theta_{111}-\theta_{0}\theta_{01}\theta_{011}}{\theta_{1}\theta_{11}(1-\theta_{111})-\theta_{0}\theta_{01}(1-\theta_{011})}\right\},$
$\operatorname{logit}(\mu_{0})$ $=\log\left\{\frac{(1-\theta_{1})\theta_{10}\theta_{101}-(1-\theta_{0})\theta_{00}\theta_{001}}{(1-\theta_{1})\theta_{10}(1-\theta_{101})-(1-\theta_{0})\theta_{00}(1-\theta_{001})}\right\}.$
Plugging $\widehat{\theta}$ into the expression of $\tau^{*}$, denoted by $\widehat{\tau}^{*}$, it is straightforward to obtain a similar result with Theorem 2. We can use the delta method to obtain the asymptotic variance for the estimator of $\operatorname{logit}(\mu_{d})$ and construct the confidence interval. The we can use the delta method again to obtain the asymptotic variance for the estimator of $\mu_{d}$ and the principal average causal effect $\tau$. Using the method of variance estimates recovery or delta method (Donner and Zou, , 2011), we can get an asymmetric confidence interval for $\tau$ which usually has better coverage rates in finite sample.
There is another equivalent approach to establish the asymptotic property for the logit transformed mean outcomes. Let $p_{zdy}=P(Z=z,D=d,S=1,Y=y)$, $\widehat{p}_{zdy}=n^{-1}\sum_{i=1}^{n}1(Z_{i}=z,D_{i}=d,S_{i}=1,Y_{i}=y)$ for $z,d,y=0,1$, and use dot $\cdot$” to represent summation with respect to its/their corresponding argument(s).
Under the identification assumptions, the point estimates of the average potential outcomes in the survived compliers are given by
$\widehat{\mu}_{1}=\frac{\frac{\widehat{p}_{111}}{r}-\frac{\widehat{p}_{011}}{1-r}}{\frac{\widehat{p}_{11\cdot}}{r}-\frac{\widehat{p}_{01\cdot}}{1-r}},\quad\widehat{\mu}_{0}=\frac{\frac{\widehat{p}_{101}}{r}-\frac{\widehat{p}_{001}}{1-r}}{\frac{\widehat{p}_{10\cdot}}{r}-\frac{\widehat{p}_{00\cdot}}{1-r}},$
or by logit transformation,
$\operatorname{logit}(\widehat{\mu}_{1})$ $=\log\left(\frac{\widehat{p}_{111}}{r}-\frac{\widehat{p}_{011}}{1-r}\right)-\log\left(\frac{\widehat{p}_{110}}{r}-\frac{\widehat{p}_{010}}{1-r}\right),$
$\operatorname{logit}(\widehat{\mu}_{0})$ $=\log\left(\frac{\widehat{p}_{101}}{r}-\frac{\widehat{p}_{001}}{1-r}\right)-\log\left(\frac{\widehat{p}_{100}}{r}-\frac{\widehat{p}_{000}}{1-r}\right).$
Theorem 3
Under Assumptions 15, suppose there is no missingness. Then
$n^{1/2}\left\{\operatorname{logit}(\widehat{\mu}_{1})-\operatorname{logit}(\mu_{1})\right\}\rightarrow_{d}N(0,V_{1}),\quad n^{1/2}\left\{\operatorname{logit}(\widehat{\mu}_{0})-\operatorname{logit}(\mu_{0})\right\}\rightarrow_{d}N(0,V_{0}),$
$n^{1/2}\left\{\tau^{*}-\tau\right\}\rightarrow_{d}N(0,V_{1}+V_{0}-2V_{10}),$
$n^{1/2}\left\{\widehat{\mu}_{1}-\mu_{1}\right\}\rightarrow_{d}N(0,\mu_{1}^{2}(1-\mu_{1})^{2}V_{1}),\quad n^{1/2}\left\{\widehat{\mu}_{0}-\mu_{0}\right\}\rightarrow_{d}N(0,\mu_{0}^{2}(1-\mu_{0})^{2}V_{0}),$
$n^{1/2}\left\{\widehat{\tau}_{0}-\tau_{0}\right\}\rightarrow_{d}N(0,\mu_{1}^{2}(1-\mu_{1})^{2}V_{1}+\mu_{0}^{2}(1-\mu_{0})^{2}V_{0}-2\mu_{1}\mu_{0}(1-\mu_{1})(1-\mu_{0})V_{10}),$
where
$V_{1}$ $=\frac{\frac{p_{111}(1-2p_{111})}{r^{2}}+\frac{p_{011}}{(1-r)^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})^{2}}+\frac{\frac{p_{110}(1-2p_{110})}{r^{2}}+\frac{p_{010}}{(1-r)^{2}}}{(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})^{2}}+\frac{\frac{4p_{111}p_{110}}{r^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})},$
$V_{0}$ $=\frac{\frac{p_{101}(1-2p_{101})}{r^{2}}+\frac{p_{001}}{(1-r)^{2}}}{(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})^{2}}+\frac{\frac{p_{100}(1-2p_{100})}{r^{2}}+\frac{p_{000}}{(1-r)^{2}}}{(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})^{2}}+\frac{\frac{4p_{101}p_{110}}{r^{2}}}{(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})},$
$V_{10}$ $=-\frac{\frac{2(p_{101}p_{000}-p_{100}p_{001})(p_{111}p_{010}-p_{110}p_{011})}{r^{2}(1-r)^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})}.$
By logit transformation, the $(1-\alpha)100\%$ asymmetric confidence intervals for $\mu_{1}$ and $\mu_{0}$ can be constructed as
$(L_{1},U_{1})$ $=\operatorname{expit}\left\{\operatorname{logit}(\widehat{\mu}_{1})\pm z_{1-\alpha/2}(V_{1}/n)^{1/2}\right\},$
$(L_{0},U_{0})$ $=\operatorname{expit}\left\{\operatorname{logit}(\widehat{\mu}_{0})\pm z_{1-\alpha/2}(V_{0}/n)^{1/2}\right\}$
respectively, where $\operatorname{expit}(x)=\exp(x)/\{1+\exp(x)\}$ and $z_{1-\alpha/2}$ is the upper $\alpha/2$ quantile of standard normal distribution. The $(1-\alpha)100\%$ asymmetric confidence interval for $\tau$ can be constructed as $(L,U)$, where
$L=\widehat{\tau}-\left\{(L_{1}-\widehat{\mu}_{1})^{2}-(U_{0}-\widehat{\mu}_{0})^{2}-2\widehat{\mu}_{0}\widehat{\mu}_{1}(1-\widehat{\mu}_{1})(1-\widehat{\mu}_{0})(L_{1}-\widehat{\mu}_{1})(U_{0}-\widehat{\mu}_{0})\right\}^{1/2},$
$U=\widehat{\tau}+\left\{(U_{1}-\widehat{\mu}_{1})^{2}-(L_{0}-\widehat{\mu}_{0})^{2}-2\widehat{\mu}_{0}\widehat{\mu}_{1}(1-\widehat{\mu}_{1})(1-\widehat{\mu}_{0})(U_{1}-\widehat{\mu}_{1})(L_{0}-\widehat{\mu}_{0})\right\}^{1/2}.$
For continuous $Y$, one can define a new outcome $Y_{c}=1_{Y>c}$ given any arbitrary threshold $c\in\mathbb{R}$. In this manner, one can obtain the estimators for the cumulative distribution functions of $Y(1)$ and $Y(0)$ in the $cl$ stratum with pointwise $(1-\alpha)100\%$ confidence intervals by varying the threshold $c$ from $\infty$- ∞ to $\infty$.
Although it is possible to derive the asymptotic properties with the existence of missingness, the expression of the asymptotic variance would be too complicated and it would be inconvenient for practical use. Therefore, the discussion with missing data is omitted here. We recommend practitioners adopt off-the-shelf tools like multiple imputation to disentangle missing data. Analyses can be performed for each completed dataset, and inference is trivial to combine these results using Rubin’s rule (Rubin, , 2018).
@@ 4 Simulation
In this section, we conduct a simulation study to assess the performance of the proposed method. Suppose that there are $n\in\{500,2000,8000\}$ independent individuals in the sample. Each individual is assigned to the treated group or control group with equal probability $r=0.5$. The compliance type is generated as
@@
So the proportion of always-takers, compliers and never-takers is 0.3, 0.28 and 0.42, respectively. The survival type is generated as
@@
To generate the outcome, we first assume that the distribution of potential outcomes are homogeneous among compliance types (Case 1). That is, $Y(d)$ does not depend on $(D(1),D(0))$. As long as $S(d)=1$, we generate
@@
Next, we consider another scenario that the distribution of potential outcomes are heterogeneous among compliance types (Case 2). For compliers,
@@
The observed variables under consistency
@@
We compare two estimators. The first estimator is the proposed one (denoted by PACE), and the standard error of the treatment effect estimator can be calculated by Theorem 2. The second estimator is the two-stage least square estimator (denoted by 2SLS), where we restrict the second-stage regression in the sample of observed survivors $\{S=1\}$. Table 2 shows the bias, standard deviation (SD), standard error (SE) and coverage rate of the nominal 95% confidence interval. The proposed method PACE has ignorable bias in Cases 1 and 2 where all assumptions hold. When principal ignorability is violated, the proposed method is biased. The 2SLS estimator is biased in all settings because it cannot deal with the cross-world reliance of $S(d)$ on $(D(1),D(0))$. With the sample size being larger, the standard deviation and standard error get smaller.
Table 2: The bias, standard deviation (SD), standard error (SE) and coverage rate of the nominal 95% confidence interval.
@@
500 Bias 0.004 0.188 -0.069 0.115 0.141 0.296 0.068 0.223
SD 0.393 0.490 0.504 0.621 0.433 0.505 0.568 0.642
SE 0.425 0.502 0.608 0.658 0.475 0.521 0.693 0.666
Coverage 0.982 0.945 0.972 0.958 0.960 0.925 0.941 0.940
2000 Bias 0.003 0.183 -0.012 0.136 0.148 0.285 0.133 0.238
SD 0.179 0.230 0.216 0.293 0.198 0.240 0.242 0.303
SE 0.184 0.244 0.255 0.319 0.203 0.253 0.288 0.323
Coverage 0.964 0.895 0.984 0.944 0.884 0.796 0.924 0.893
8000 Bias -0.003 0.180 -0.005 0.142 0.146 0.283 0.144 0.245
SD 0.090 0.118 0.111 0.151 0.101 0.125 0.126 0.160
SE 0.089 0.122 0.124 0.159 0.099 0.126 0.140 0.161
Coverage 0.943 0.701 0.968 0.858 0.670 0.379 0.810 0.659
5 Application to Job Corps data
The U.S. Job Corps (JC) training program was a randomized education and job training program for disadvantaged youths. The study collected baseline covariates including gender, age, race, education status and family information. There are 9240 complete cases, among which 5577 were assigned to Job Corps and 3663 out of Job Corps. In the first year after assignment, 84.6% enrolled in education and/or vocational training in the treated group, and 50.6% enrolled in education and/or vocational training in the control group. In the second year after assignment, 52.3% enrolled in education and/or vocational training in the treated group, and 42.0% enrolled in education and/or vocational training in the control group. In our analysis, we regard the individual as receiving the training if he/she has ever enrolled in the training when analyzing the treatment effects later than the second year after assignment. So 88.8% in the treated group actually received the treatment and 60.6% in the control group actually received the treatment. We infer that the proportion of always-takers, compliers and never-takers is 0.606, 0.282 and 0.112, respectively.
In one year after assignment, 49.4% have ever worked in the treated group, and 57.1% have ever worked in the control group. In the second year after assignment, the proportion of employment is 75.2% in the treated group and 75.3% in the control group. In the third year after assignment, the proportion of employment increases to 83.5% in the treated group and 81.4% in the control group. In the fourth year after assignment, the proportion of employment increases to 83.7% in the treated group and 81.3% in the control group. We see that the employment status becomes stable in three years after assignment. We take the logarithm for the weekly earnings. There is a trend that the earnings grow higher with time going on. Table 3 lists the proportion of employment and average earnings in each level of the assigned treatment $Z$ and received treatment $D$.

### v2

@@ 2.2 Identification of the PACE
$P(S(0)=1\mid c)$ =$=\frac{\{p_{c}+p_{n}\}E\{S\mid Z=0,D=0\}-p_{n}E\{S\mid Z=1,D=0\}}{p_{c}}.$
The third-stage regression links the outcomes with principal strata. Note that $YS$ is always well-defined,
@@
Afterward, the identifiability of PACE follows immediately.
@@
In particular, we have the following results as corollaries of Theorem 1 when the identification assumptions hold.
@@
In the presence of missingness, we can use standard multiple imputation methods to impute the missing values and use the formula in Remark 3 to estimate the principal average causal effect (Harel and Zhou, , 2007; Kenward and Carpenter, , 2007; Rubin, , 2018).
It is noted that the proportions of principal strata $cl$, $cp$, $ch$, and $cd$ defined by the compliance type and survival type are not identifiable without additional assumptions. To identify the proportion of survived compliers, a typical assumption is the individual-level monotonicity on $S$, i.e., $S(1)\geq S(0)$ almost surely, under which
@@
Whether this assumption is plausible depends on specific scientific questions. For example, this monotonicity does not hold for the Job Corps data because joining the training may raise psychological expectations about wages (Frumento et al., , 2012). Still, it may be reasonable in some medical applications where the active treatment is beneficial for survival (Ding et al., , 2011; Wang et al., , 2017). The non-identifiability of principal strata proportions brings difficulty aggregating the conditional causal effect at each covariate level to the average causal effect in the overall population. More details about identification are provided in the Supplementary Material.
@@ 3 Estimation and inference
Consider a completely randomized experiment of size $n$ with a possibly unknown treatment probability $P(Z=1)=1-P(Z=0)=r$. Suppose there is no missingness; otherwise, one can use multiple imputations or other methods to generate a completed dataset first. The asymptotic variance of the estimator can be calculated by Rubin’s rule from completed datasets (Rubin and Schenker, , 1986; White et al., , 2011; Rubin, , 2018).
@@
$\theta=(r,\theta_{1},\theta_{0},\theta_{11},\theta_{10},\theta_{01},\theta_{00},\theta_{111},\theta_{101},\theta_{011},\theta_{001})^{{\mathrm{\scriptscriptstyle T}}}.$
The likelihood for $\theta$ is
@@
The maximized likelihood estimate of $\theta$ is given by
$\widehat{\theta}=\left(\frac{n_{1\cdot\cdot}}{n},\frac{n_{11\cdot}}{n_{1\cdot\cdot}},\frac{n_{01\cdot}}{n_{0\cdot\cdot}},\frac{n_{111}}{n_{11\cdot}},\frac{n_{101}}{n_{10\cdot}},\frac{n_{011}}{n_{01\cdot}},\frac{n_{001}}{n_{00\cdot}},\widehat{\theta}_{111},\widehat{\theta}_{101},\widehat{\theta}_{011},\widehat{\theta}_{001}\right)^{{\mathrm{\scriptscriptstyle T}}},$
@@
$\operatorname{cov}(\widehat{\theta})$ $=\operatorname{diag}\bigg{\{}\frac{r(1-r)}{n},\frac{\theta_{1}(1-\theta_{1})}{nr},\frac{\theta_{0}(1-\theta_{0})}{n(1-r)},\frac{\theta_{11}(1-\theta_{11})}{nr\theta_{1}},\frac{\theta_{10}(1-\theta_{10})}{nr(1-\theta_{1})},$
$\qquad\qquad\frac{\theta_{01}(1-\theta_{01})}{n(1-r)\theta_{0}},\frac{\theta_{00}(1-\theta_{00})}{n(1-r)(1-\theta_{0})},\operatorname{var}(\widehat{\theta}_{111}),\operatorname{var}(\widehat{\theta}_{101}),\operatorname{var}(\widehat{\theta}_{011}),\operatorname{var}(\widehat{\theta}_{001})\bigg{\}}.$
If we use the sample average $\widehat{\theta}_{zd1}$ of the outcomes in $\{Z=z,D=d,S=1\}$ to estimate the conditional mean outcome $b(\theta_{zd1})$, then $\operatorname{var}(\widehat{\theta}_{zd1})$ is the variance of the observed outcomes in $\{Z=z,D=d,S=1\}$ divided by the group size $n_{zd1}$.
Since we can express the mean potential outcome $\mu_{d}=E\{Y(d)\mid G=cl\}$ as
@@
the causal paratemers can be estimated by plugging $\widehat{\theta}$ into $\mu_{d}$, denoted by $\widehat{\mu}_{d}$, for $d=0,1$. The principal average causal effect is estimated by $\widehat{\tau}=\widehat{\mu}_{1}-\widehat{\mu}_{0}$. The asymptotic properties of the plug-in estimators can be established by the sandwich estimator, as shown in the following theorem.
@@
When the outcome $Y$ is binary, the normal approximation may not have a satisfactory finite-sample performance, since the distribution of the estimator $\widehat{\mu}_{d}$ is skewed. To respect the fact that $\mu_{d}$ must lie in $[0,1]$, we may as well adopt $\operatorname{logit}(\mu_{d})=\log\{\mu_{d}/(1-\mu_{d})\}$ as the target estimand, which can range on the whole real line. The contrast $\tau^{*}=\operatorname{logit}(\mu_{1})-\operatorname{logit}(\mu_{0})$ reflects the log odds ratio between treatment and control for survived compliers. However, the log odds ratio does not enjoy collapsibility. The asymptotic properties for the maximized likelihood estimator of $\tau^{*}$ are provided in the Supplementary Material. In the presence of missingness, analyses can be performed for each completed dataset, and inference is trivial to combine these results using Rubin’s rule (Rubin, , 2018).
@@ 4 Simulation
In this section, we conduct a simulation study to assess the performance of the proposed method. Suppose there are $n\in\{500,2000,8000\}$ independent individuals in the sample. Each individual is assigned to the treated or control group with equal probability $r=0.5$. The compliance type is generated as
@@
So the proportion of always-takers, compliers, and never-takers is 0.3, 0.28, and 0.42, respectively. The survival type is generated as
@@
The compliance type $(D(1),D(0))$ acts as the unmeasured confounder $U$, which can affect both $D$ and $S$.
To generate the outcome, we consider two settings. We first assume that the distribution of potential outcomes is homogeneous among compliance types (Case 1). That is, $Y(d)$ does not depend on $(D(1),D(0))$. As long as $S(d)=1$, we generate
@@
Next, we consider another scenario where the distribution of potential outcomes is heterogeneous among compliance types (Case 2). For compliers,
@@
The observed variables under consistency are
@@
We compare two estimators. The first estimator is the proposed one (denoted by PACE), and the standard error of the treatment effect estimator can be calculated by Theorem 2. The second estimator is the two-stage least squares estimator (denoted by 2SLS), where we restrict the second-stage regression to the sample of observed survivors $\{S=1\}$. Table 2 shows the bias, standard deviation (SD), standard error (SE), and coverage percentage (CP) of the nominal 95% confidence interval in 10000 repeatedly generated datasets. The proposed method, PACE, has a negligible bias in Cases 1 and 2, where all assumptions hold. When principal ignorability is violated in Cases 3 and 4, the proposed method is biased, but the bias is smaller than that of 2SLS. The 2SLS estimator is biased in all settings because it cannot deal with the cross-world reliance of $S(d)$ on $(D(1),D(0))$. The confidence interval constructed by the asymptotic formula has a coverage percentage close to the nominal level. In all settings, the proposed method has a samller variance the 2SLS, indicating the efficiency of the proposed method. With the larger sample size, the standard deviation and standard error get smaller.
Table 2: The bias, standard deviation (SD), standard error (SE), and coverage percentage (CP) of the nominal 95% confidence interval in the simulation study
@@
500 Bias 0.006 0.191 -0.064 0.124 0.142 0.293 0.072 0.225
SD 0.394 0.475 0.497 0.600 0.438 0.498 0.572 0.628
SE 0.402 0.499 0.490 0.654 0.444 0.518 0.551 0.662
CP 0.969 0.948 0.966 0.963 0.948 0.925 0.933 0.944
2000 Bias -0.000 0.180 -0.014 0.135 0.146 0.283 0.132 0.238
SD 0.177 0.224 0.215 0.287 0.197 0.239 0.242 0.302
SE 0.182 0.244 0.216 0.320 0.200 0.254 0.240 0.324
CP 0.958 0.905 0.954 0.947 0.885 0.812 0.890 0.895
8000 Bias 0.001 0.183 -0.002 0.144 0.150 0.287 0.147 0.248
SD 0.088 0.114 0.106 0.145 0.098 0.120 0.119 0.151
SE 0.089 0.122 0.105 0.159 0.098 0.126 0.117 0.161
CP 0.948 0.689 0.948 0.875 0.658 0.369 0.733 0.668
In summary, the proposed method can consistently estimate the principal average causal effect. When the core identification assumption is violated, it is still more robust than the 2SLS.
5 Application to Job Corps study
The U.S. Job Corps training program was a randomized education and job training program for disadvantaged youths. The study collected baseline covariates including gender, age, race, education status, and family information. There are 9240 complete cases, among which 5577 were assigned to Job Corps and 3663 were not assigned to Job Corps. In the first year after assignment, 84.6% enrolled in education and/or vocational training in the treated group, and 50.6% enrolled in education and/or vocational training in the control group. In the second year after the assignment, 52.3% enrolled in education and/or vocational training in the treated group, and 42.0% enrolled in education and/or vocational training in the control group. In our analysis, we regard the individual as receiving the training if he/she has ever enrolled in the training when analyzing the treatment effects later than the second year after assignment. So, 88.8% in the treated group actually received the treatment, and 60.6% in the control group actually received the treatment. We infer that the proportion of always-takers, compliers, and never-takers is 0.606, 0.282, and 0.112, respectively.
In one year after the assignment, 49.4% have ever worked in the treated group, and 57.1% have ever worked in the control group. In the second year after the assignment, the proportion of employment is 75.2% in the treated group and 75.3% in the control group. In the third year after assignment, the proportion of employment increases to 83.5% in the treated group and 81.4% in the control group. In the fourth year after the assignment, the proportion of employment increases to 83.7% in the treated group and 81.3% in the control group. We see that the employment status becomes stable in three years after the assignment. We take the logarithm of the weekly earnings. There is a trend for earnings to grow higher over time. Table 3 lists the proportion of employment and average earnings in each level of the assigned treatment $Z$ and received treatment $D$.